OpenStudy (anonymous):

Use a sum or difference identity to determine the exact value of cos(5π/12)

OpenStudy (mertsj):

Have you thought of 9/12-4/12?

OpenStudy (mertsj):

Or maybe 8/12-3/12?

OpenStudy (anonymous):

How would you get the \[\pi\] in your answer, I can see how the those two things you mentioned would give me 5/12.

OpenStudy (mertsj):

\[\frac{5\pi}{12}=\frac{8\pi}{12}-\frac{3\pi}{12}=\frac{2\pi}{3}-\frac{\pi}{4}\]

OpenStudy (mertsj):

\[\cos \frac{5\pi}{12}=\cos (\frac{2\pi}{3}-\frac{\pi}{4})=\cos \frac{2\pi}{3}\cos \frac{\pi}{4}+\sin \frac{2\pi}{3}\sin \frac {\pi}{4}\]

OpenStudy (anonymous):

Thank you very much, I will use the sum and difference idenities to find the exact value of \[\cos (2\pi/3-\pi/4)\]

OpenStudy (mertsj):

yw